please in a clear answer find me those:
the yield stress
the ultimate tensile stress(uts)
breaking strain
50.30 11.15 6.00 26.00 R1.00 0 11.00 R1.50 04.00 07.14
Load (N) Tool Steel 25000 Break 20000 15000 10000 5000 000 3.0 40 00 20 Machine Extension (mm)
Answer #1

O For the given Speciomen A = Id = I tu - TxY DA= D/A = 12.56 mm . 14500 Yield Stress = 1393.3 Ty 12.56 1. Ty = 1393.3.1/mm =Lo= gauge length = 26mm Breaking Strain = 3.8. AL Lo 0.1467 26 F. Breaking Strain = 0.146 es If - final length LotAL 26+3.8 =